An isosceles triangle has 2 equal sides and the other side is 5 less than the sum of those 2 sides. The perimeter around the triangle is 23. What is the length of one of the equal sides.

Answer by peekaboo0501(13)

23= 3x+x-5
23+5=4x
7 = x - the leghth of two equal side
14 - 5 = 9 - the length of the other side